OpenStudy (anonymous):

Simplify by factoring. sqrt(20x^2y)

jhonyy9 (jhonyy9):

so how can you factorizing 20 ?

jhonyy9 (jhonyy9):

using a number like perfect square

jhonyy9 (jhonyy9):

so will be sqrt(4*5*x^2 *y) =

jhonyy9 (jhonyy9):

= sqrt(4x^2) * sqrt(5y) = 2xsqrt(5y) hope you can understanding it sure easy

OpenStudy (anonymous):

I understand it now, thank you.

jhonyy9 (jhonyy9):

was my pleasure you are welcom good luck bye